OpenStudy (anonymous):

10x^-11x-6=0 solve using appropriate technique????

OpenStudy (lgbasallote):

1) what's 10x^ 2) what appropriate technique do you mean?

OpenStudy (anonymous):

10x^2-15x+4x-6=0 5x(2x-3)+2(2x-3)=0 (2x-30(5x+2)=0 x=3/2 x=-2/5